What is the area of a rectangle with length of 16 inscribed in a circle with a radius of 10
Answer by prateekagrawal(56)

here... diameter of circle= diagonal of rectangle.
so,
diameter= 2*10= 20cm.
so, diagonal of rectangle= 20cm.
(diagonal)^2 = length^2 + width^2
so,
20^2 = 16^2 +w^2
144= w^2
w=12 cm.
so,
area of rectangle = length* width = 16*12 = 192 cm^2.